# AI Copilots for Reproducibility in Science: A Case Study

### v3

@@ Conclusion
As shown in the study by Gundersen et al. (2025), current barriers to reproducibility often stem from missing or unclear information [9]. In our evaluation, we showed that in a case that was documented in Gundersen et al., our AI system reduced the reproduction time from more than 30 hours to approximately 1 hour. This efficiency gain illustrates the value of AI-generated proactive guidance tailored to both authors and readers.
Importantly, the success of the system stems from its ability to surface specific, high-impact shortcomings - missing hyperparameters, inaccessible datasets, ambiguous code - and to communicate those deficiencies in ways that are actionable for authors. It also provides reproducibility scaffolds for readers, such as structured Jupyter Notebooks.

### v4

@@ Conclusion
As shown in the study by gundersen2025unreasonableeffectiveness, current barriers to reproducibility often stem from missing or unclear information. In our evaluation, we showed that in a case documented in Gundersen et al., our AI system reduced the reproduction time from more than 30 hours to approximately 1 hour. This efficiency gain illustrates the value of AI-generated proactive guidance tailored to both authors and readers.
Importantly, the success of the system stems from its ability to surface specific, high-impact shortcomings - missing values for hyperparameters, inaccessible datasets, ambiguous code - and to communicate those deficiencies in ways that are actionable for authors. It also provides reproducibility scaffolds for readers, such as structured Jupyter Notebooks.